OpenStudy (anonymous):

using the equation p-7=5/57(h-35) what is the predicted number of pounds to come from an 18 food tree? a 20 foot tree?

OpenStudy (anonymous):

First solve for p by adding 7 to both sides: p = 5/57(h-35) + 7. Then simplify the equation: Then plug in 18 for h and do the math. Then plug in 20. Here's how to do 18: p = 5/57(18-35) + 7 p = 5/57(-17) + 7 p = -85/57 + 7 p = -85/57 + 399/57 p = 314/57 p = 5.51